### v1

# Graph Network Simulator MGN MeshGraphNet

GNS
Graph Network Simulator
MGN
MeshGraphNet
GNN
Graph Neural Network
HGNN
Hierarchical Graph Neural Network
MPN
Message Passing Network
MLP
Multilayer Perceptron
CNN
Convolutional Neural Network
MSE
Mean Squared Error
PDE
Partial Differential Equation
GMM
Gaussian Mixture Model
FEM
Finite Element Method
FVM
Finite Volume Method
FDM
Finite Difference Method
ML
Machine Learning
AMG
Algebraic multigrid
DDPM
Denoising Diffusion Probalistic Model
UMGN
U-hierarchical MeshGraphNets
AMPN
Algebraic-hierarchical Message Passing Network
U-Net
U-Net
ROBIN
Rolling Diffusion-Batched Inference Network
ROBI
Rolling Diffusion-Batched Inference
SiLU
Sigmoid Linear Unit
PDE-Refiner
PDE-Refiner
HCMT
Hierarchical Contact Mesh Transforme
AMR
Adaptive Mesh Refinement
BSDL
Bi-stride-coarsening and Delaunay-remeshing
Intra-MP-Stack
Intra-Level-Message Passing Stack
Inter-MP-Stack
Inter-Level-Message Passing Stack
CMT
Contact Mesh Transformer
HMT
Hierarchical Mesh Transformer
Diffusion-Based Hierarchical Graph Neural Networks for Simulating Nonlinear Solid Mechanics
Tobias Würth1   Niklas Freymuth2  Gerhard Neumann2  Luise Kärger1
1Institute of Vehicle System Technology, Karlsruhe Institute of Technology, Karlsruhe
2Autonomous Learning Robots, Karlsruhe Institute of Technology, Karlsruhe
correspondence to tobias.wuerth@kit.edu
Abstract
Graph-based learned simulators have emerged as a promising approach for simulating physical systems on unstructured meshes, offering speed and generalization across diverse geometries. However, they often struggle with capturing global phenomena, such as bending or long-range correlations, and suffer from error accumulation over long rollouts due to their reliance on local message passing and direct next-step prediction. We address these limitations by introducing the Rolling Diffusion-Batched Inference Network (ROBIN), a novel learned simulator that integrates two key innovations: (i) Rolling Diffusion, a parallelized inference scheme that amortizes the cost of diffusion-based refinement across physical time steps by overlapping denoising steps across a temporal window. (ii) A Hierarchical Graph Neural Network built on algebraic multigrid coarsening, enabling multiscale message passing across different mesh resolutions. This architecture, implemented via Algebraic-hierarchical Message Passing Networks, captures both fine-scale local dynamics and global structural effects critical for phenomena like beam bending or multi-body contact. We validate ROBIN on challenging $2$D and $3$D solid mechanics benchmarks involving geometric, material, and contact nonlinearities. ROBIN achieves state-of-the-art accuracy on all tasks, substantially outperforming existing next-step learned simulators while reducing inference time by up to an order of magnitude compared to standard diffusion simulators.
1 Introduction
Physical simulations enable many engineering and scientific fields to gain quick insights into complex systems or to evaluate design decisions. Conventional simulations model the physical system using Partial Differential Equations. Usually, the PDE is discretized by numerical methods, such as the  Finite Element Method (FEM[1], the  Finite Volume Method (FVM[2] or the  Finite Difference Method (FDM[3]. This process reduces the need for cumbersome, resource-intensive real-world experiments. Recent research aims to speed up simulation with  Machine Learning (ML)-based models [4, 5]. These learned simulators promise to allow researchers and practitioners to evaluate large amounts of virtual, simulated experiments. These simulations in turn unlock applications in engineering design and manufacturing optimization [6, 7, 8].
This work aims to improve learned simulators, focusing on simulations from nonlinear solid mechanics as a representative class of examples. We combine recent image-based Denoising Diffusion Probalistic Models [9, 10, 11] with  Hierarchical Graph Neural Network (HGNN[12, 13, 14]. While diffusion has shown promising results on images [15, 16, 17], audio [18, 19] and even policy learning for robotics [20, 21], it suffers from cost-intensive inference due to its iterative denoising procedure. We alleviate this high inference cost on time-dependent domains with   Rolling Diffusion-Batched Inference (ROBI), a novel scheduling scheme that batches denoising steps of consecutive time steps. ROBI already starts denoising future prediction steps by using partially refined previous steps. This process reduces the number of model evaluations to the number of time steps and preserves the time-shift equivarinace of Markovian systems. ROBI only affects the inference processes, allowing us to utilize conventional, parallelized DDPM training.
Figure 1: Overview of ROBINROBIN coarses the fine mesh multiple times with AMG to create a graph hierarchy. ROBIN predicts the denoising velocity batch $\mathbf{v}_{\theta,k_{\text{B}}}^{t_{\text{B}}}$ using AMPNs, given a batch of noisy samples $\Delta{\mathbf{u}}^{t_{\text{B}}}_{k_{\text{B}}}$ and clean samples $\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0|k_{\text{B}}+1}$ as input. The prediction $\mathbf{v}_{\theta,k_{\text{B}}}^{t_{\text{B}}}$ is used for a denoising step of the noisy sample $\Delta{\mathbf{u}}^{t_{\text{B}}}_{k_{\text{B}}-1}$ and a state update $\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0|k_{\text{B}}}$.
We combine ROBI with a diffusion HGNN to form the  Rolling Diffusion-Batched Inference Network (ROBIN), which significantly accelerates simulation while improving predictive accuracy. ROBIN constitutes the first diffusion-based refiner for simulating physical dynamics on unstructured meshes, surpassing the current accuracy ceiling of hierarchical Graph Neural Networks. We train ROBINs on three challenging $2$D and $3$D solid mechanics datasets involving geometric, material, and contact nonlinearities. Across all datasets, ROBIN significantly improves over state-of the-art mesh-based simulators [5, 22] in terms of predictive accuracy. We further find that our proposed inference method, ROBI, speeds up diffusion-based inference for learned simulations by up to an order of magnitude while maintaining accuracy. [^1] Code and datasets will be open-sourced upon acceptance.
To summarize, our contributions we i) propose ROBI, a novel inference scheduling scheme for diffusion-based simulators that amortizes denoising across time steps, reducing inference to a single model evaluation per step while preserving time-shift equivariance; ii) introduce ROBIN, a diffusion-based HGNN for nonlinear solid mechanics that combines multiscale message passing with ROBI to provide fast, accurate diffusion-based simulations; iii) demonstrate state-of-the-art performance on challenging $2$D and $3$D benchmarks, outperforming existing simulators in both accuracy and runtime.
2 Related Work
Simulating Complex Physics. Simulating complex physical systems often requires numerical solvers, such as the FEM [23, 24, 1], the FVM [2], or the FDM [3]. While accurate, numerical solvers scale poorly with problem complexity, often requiring multiple hours of even days for a single rollout on a modern workstation. Recent work shows that ML-based models are able to learn such numerical simulations from data [25, 26, 4, 5, 27, 9, 22, 10]. ML-based models provide speed-ups of one to two orders of magnitude while being fully differentiable, which accelerates downstream applications such as design [6] or manufacturing process optimization [7, 8]. Many learned simulators operate autoregressively. They mimic numerical solvers by using their own predictions to estimate the residual between successive time steps [4, 5, 28]. Similarly, Neural ODEs [29, 30] predict time derivatives and advance solutions via numerical integration. In contrast to these supervised approaches, Physics-Informed Neural Networks [31] directly operate on a PDEs loss function to train Multilayer Perceptrons [32, 33, 8]Convolutional Neural Networks [34, 35] or GNN [36, 37]. Finally, Neural Operators aim to learn mesh-independent solution operators [38, 39, 40, 27, 41]. ROBIN is an autoregressive learned simulator that replaces direct next-step prediction with multiple denoising diffusion steps, leveraging generative inference to improve prediction accuracy.
Learning Mesh-based Simulations with Graph Neural Network (GNN). Pfaff et al. [5] introduced MeshGraphNets, a Message Passing Network (MPN)-based simulator that encodes simulation states as graphs using mesh connectivity and physical proximity. While accurate for small problems, MeshGraphNet (MGN) does not scale well, as its receptive field is limited by the number of message-passing steps in the MPN. To address this issue, recent work expands the receptive field via global attention [42, 43, 44] or hierarchical mesh representations using Graph Convolutional Networks [45] and MPNs [46, 13]. Extensions further improve efficiency and accuracy through rotation equivariance [14], hierarchical edge design [47], bi-stride pooling [48], and attention mechanisms acting on edges and across hierarchies [22]. [49] leverages  Adaptive Mesh Refinement (AMR) to create mesh hierarchies for multi-scale GNNs. Existing methods are generally trained using a next-step Mean Squared Error (MSE) loss, which favors learning lower solutions frequencies at the cost of accuracy in higher frequency bands that have less impact on the loss [9]. However, autoregressive models trained with a MSE loss often overlook low-amplitude frequencies [9]. Our approach is orthogonal, coupling hierarchical GNN with denoising diffusion models. This approach allows to take advantage of the large receptive field of multi-scale GNNs while pushing accuracy toward diffusion limits.
Diffusion-based Simulations. Diffusion models have been applied to physics-informed image super-resolution [35], flow field reconstruction [50], and steady-state flow generation on grids using CNNs [51], and more recently to meshes with hierarchical GNNs [52]. These models, however, operate on isolated frames and do not capture time-dependent dynamics. In contrast, our model predicts deterministic physical evolution rather than equilibrium samples via autoregressive rollouts.
While next-step simulators trained with MSE loss capture high-amplitude, low-frequency components, they often miss low-amplitude components [9]. PDE-Refiners (PDE-Refiners) addresses this via iterative refinement, improving long-horizon accuracy of grid-based CNN simulators. We extend this idea to unstructured meshes by combining algebraic mesh coarsening [53, 54] with hierarchical GNNs [52, 22] with shared layers. We further introduce a time-parallel denoising scheme at inference, removing the speed bottleneck while maintaining accuracy. Unlike diffusion-based CNN simulators that require $K$ model evaluations per physical step $T$ [11, 55, 9], our method requires only a single hierarchical GNN call per step post warm-up, reducing inference from $\mathcal{O}(KT)$ to $\mathcal{O}(T)$. Compared to video-based approaches [10, 56], which need to jointly process $N$ steps and learn a time-dependent denoiser with high memory cost, our model, ROBIN, leverages time-translation invariance to train a time-independent denoiser with only one time step in memory. As such, ROBIN can be applied autoregressively and can freely interpolate between fully parallel denoising and memory-efficient one-step denoising at inference time. It also predicts state residuals instead of states, significantly improving long-horizon rollout fidelity.
3   Rolling Diffusion-Batched Inference Network (ROBIN)
Graph Network Simulators (GNSs) for Mesh-based Simulations. We consider solving PDEs for physical quantities $\mathbf{u}(\mathbf{x},t)$ that change over time $t\in[0,T]$ and inside a time-dependent domain $\mathbf{x}(t)\in\Omega(t)$. We focus on simulations on meshes, where $\mathcal{G}=(\mathcal{V},\mathcal{E}^{\text{M}})$ denotes the mesh graph and the graph nodes $\mathcal{V}$ and the graph edges $\mathcal{E}^{\text{M}}$ correspond to mesh nodes and mesh edges. We seek solutions $\mathbf{u}_{i}(t)=\mathbf{u}(\mathbf{x}_{i},t)$ at discrete node locations $\mathbf{x}_{i}(t)\in\Omega(t)$. To obtain discretized PDEs, usually numerical methods, such as the FEM, are applied that define the discretization of spatial operators, such as gradients $\partial\mathbf{u}(\mathbf{x},t)/\partial\mathbf{x}$. Given the discretized operator $\mathcal{F}$, the PDE simplifies to an time-dependent Ordinary Differential Equation and requires solving $\partial\mathbf{u}_{i}/\partial t=\mathcal{F}(t,\mathbf{x}_{i},\mathbf{u}_{i})~{}\text{.}$ We can solve such systems using numerical time discretization schemes. In this work we use a simple Euler forward discretization $\mathbf{u}_{i}^{t{+}1}=\mathbf{u}_{i}^{t}+\Delta t~{}\mathcal{F}(t,\mathbf{x}_{i}^{t},\mathbf{u}_{i}^{t})~{}\text{,}$ and set $\Delta t=1$. We extend  PDE-Refiner (PDE-Refiner[9] to Lagrangian systems, where the domain $\Omega=\Omega^{t}$ and node locations $\mathbf{x}_{i}=\mathbf{x}_{i}^{t}\in\Omega^{t}$ evolve over time. Here, we predict the solution $\mathbf{u}_{i}^{t}$ at time step $t$ by learning to reverse a probabilistic diffusion process[57] conditioned on the solution history. The proposed methods also apply without any restriction to Eulerian systems, where the domains are time-independent.
3.1 Denoising Diffusion Probalistic Models (DDPMs) for time-dependent simulations
Given a time-dependent solution $\mathbf{u}^{t}$ from a data distribution $q(\mathbf{u})$, the forward diffusion process is modeled by a Markov chain, where Gaussian noise is added gradually to the sample $\mathbf{u}^{t}_{k}$ at each diffusion step $k$
$q(\mathbf{u}^{t}_{1:K}|\mathbf{u}^{t}_{0})=\prod_{k=1}^{K}q(\mathbf{u}^{t}_{k}|\mathbf{u}^{t}_{k-1})~{}\text{,}~{}~{}~{}~{}q(\mathbf{u}^{t}_{k}|\mathbf{u}^{t}_{k-1}):=\mathcal{N}(\mathbf{u}^{t}_{k};\sqrt{1-\beta_{k}}\mathbf{u}^{t}_{k-1},\beta_{k}\mathbf{I})~{}\text{.}$
$\mathcal{N}$ denotes the normal distribution and $\beta_{k}$ the noise specified by a variance scheduler. We learn to reverse the diffusion process, i.e.,
$p_{\theta}(\mathbf{u}^{t}_{k-1}|\mathbf{u}^{t}_{k}):=\mathcal{N}(\mathbf{u}^{t}_{k-1};\mu_{\theta}(\mathbf{u}^{t}_{k},k),\Sigma_{\theta}(\mathbf{u}^{t}_{k},k))~{}\text{,}$
where the mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$ are predicted by a learned model with parameters $\theta$. The covariance is assumed to be isotropic and given as $\sigma_{k}^{2}\mathbf{I}=(\frac{1-\bar{\alpha}_{k-1}}{1-\bar{\alpha}_{k}}\beta_{k})\mathbf{I}$ [57] with $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{k}=\prod_{i=1}^{k}\alpha_{i}$. The reverse denoising process consists of $K$ iterative diffusion steps and starts from $k=K$. To facilitate faster denoising, we follow the DDPM formulation of [9] and use an exponential $\beta_{k}$ scheduler. We train the model to predict the denoising velocity, i.e., the v-prediction target
$\mathbf{v}_{k}^{t}=\sqrt{\bar{\alpha}_{k}}\mathbf{\epsilon}^{t}-\sqrt{1-\bar{\alpha}_{k}}\mathbf{u}^{t}_{0}~{}\text{,}$ (1)
given gaussian noise $\mathbf{\epsilon}^{t}$ [58]. We predict the solution autoregressevily, conditioning the model on the last time step solution $\mathbf{u}^{t-1}_{0}$ and additional conditioning features $\mathbf{c}^{t-1}$ to predict the current denoising velocity $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k)=\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$. We define the target as $\mathbb{E}_{\mathbf{u}^{t}_{0},\mathbf{\epsilon}^{t},k}\left[\left\|\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})-\mathbf{v}_{k}^{t}\right\|^{2}\right]$, which corresponds to a MSE loss over predicted denoising velocities [9]. The first denoising step is defined such that $\bar{\alpha}_{K}\approx 0$, which simplifies the v-prediction target to $\mathbf{v}_{k}^{t}\approx-\mathbf{u}^{t}_{0}$ (cf. Equation 1). The noisy sample $\mathbf{u}^{t}_{k}\approx\mathbf{\epsilon}^{t}$ corresponds to Gaussian Noise and is uninformative. Hence, for $k=K$ the model target converges to $\left\|\mathbf{v}_{\theta}(\mathbf{u}^{t}_{K},K,\mathbf{u}^{t-1}_{0}))-\mathbf{v}_{k}^{t}\right\|^{2}\approx\left\|\mathbf{v}_{\theta}(\mathbf{\epsilon}^{t},K,\mathbf{u}^{t-1}_{0}))+\mathbf{u}^{t}_{0}\right\|^{2}$, i.e., the MSE objective of one step models. Consequently, the predicted solution after the first denoising step $k=K$ mirrors that of one-step models [9]. More importantly, due to the exponential noise scheduler, each denoising step focuses on different amplitude and frequency levels of the solution [9], with later denoising steps increasingly paying attention to higher frequencies.
Rolling Diffusion-Batched Inference (ROBI). Conventional diffusion inference requires $K$ model calls, each corresponding to a denoising step, per simulation time step [9]. Figure 2 a) shows an example. Thus, inference is roughly $K$ times slower than one-step models [5]. We propose  Rolling Diffusion-Batched Inference (ROBI) to accelerate inference in DDPM-based autoregressive simulators. Given a velocity prediction $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$ at the denoising step $k$, we reconstruct a partially refined prediction $\tilde{\mathbf{u}}^{t}_{0|k}=\sqrt{\bar{\alpha}_{k}}\mathbf{u}^{t}_{k}-\sqrt{1-\bar{\alpha}_{k}}\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$  [58].
Figure 2: Overview of a) conventional autoregressive diffusion inference and b) ROBI. a) Conventional inference denoises the entire state of a physical time step at once before it shifts to the next time step (see One Step Denosing). The Iterative Inference requires $K$ model calls per time step, where $K$ denotes the number of diffusion steps. b) ROBI parallelizes denoising steps across physical time, processing up to $K$ time steps batched, and reconstruct the physical states with the clean sample prediction subsequently (see Batched Denoising Step). This process allows Rolling-Batched Inference after the initial warm-up, reducing the number of model calls to one per time step.
As discussed in Section 3.1, early denoising prediction steps resemble one-step models, while later steps progressively refine spatial frequency bands, from coarse structures to fine details. After $m<K$ denoising steps at time $t$, the intermediate prediction $\tilde{\mathbf{u}}^{t}_{0|m}$ thus already captures low-to-mid frequency information. We assume this information is sufficient to condition the next step $t{+}1$, partiularly to predict a solution within this already refined frequency band. Using these properties, we begin denoising of time step $t{+}1$ after $m$ steps by initializing $\mathbf{u}^{t{+}1}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, allowing batched denoising of steps $t{:}t{+}1$. More generally, after $hm$ denoising steps, we initialize the time step $t+h$. We denoise a rolling prediction window of $t:t+h$ time steps in parallel. After warm up, this process yields a constant prediction window of size $h=K/m$ and reduces the number of model calls from $KT$ to $K-m+mT$ steps. We denote $m$ as denoising stride.
Figure 2 b) visualizes this for the special case $m=1$ and $K=3$ diffusion steps. Each model call moves the simulation forward in time by one step, denoising each of the $K$ partially denoised predictions in its rolling windows by one additional noise step. A denoising step of ROBI consists of two consecutive steps. First the learned model predicts the denoising velocity $\mathbf{v}_{\theta}(\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}},k_{\text{B}},\mathbf{u}^{t_{\text{B}}-1}_{0})$ of the prediction horizon batch with the time steps $t_{\text{B}}=t:t+h$ and denoising steps $k_{\text{B}}=m:m:K$. Subsequently, the denoising scheduler is used to reconstruct the clean samples $\tilde{\mathbf{u}}^{t_{b}}_{0|k_{b}}$ and to update the noisy samples $\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}-1}\sim p_{\theta}(\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}-1}|\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}})$. After each model call with a batch size of $h=3$, the prediction window shifts one physical time step forward.
For $m=K$ROBI converges to conventional autoregressive diffusion inference (cf. Figure 2 a)). Thus, the denoising stride $m$ can be considered a hyperparameter that trades off prediction accuracy and memory usage with inference speed. Notably, ROBI does not affect the training process or the model architecture. The model remains a one-step diffusion model, which preserves the time-shift equivariance of Markovian systems, fast training convergence and small GPU memory utilization. Furthermore, we only require the last time step as history, which proves to be more stable and accurate for autoregressive ML-based simulations [5, 9].
In practice, we find that denoising residual predictions $\Delta\mathbf{u}^{t}_{k}$ instead of $\mathbf{u}^{t}_{k}$ improves accuracy. Let $\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0}\approx\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0|k_{\text{B}}}$ be the predicted clean residuals with $t_{\text{B}}=t:t+h$ and denote the last fully denoised state as $\mathbf{u}^{t-1}_{0}$. We extend ROBI to update the clean state estimate $\tilde{\mathbf{u}}^{t+j}_{0}$ of the time step $t+j$, $j\in\{0,h\}$ in the prediction window using the cumulative sum $\tilde{\mathbf{u}}^{t+j}_{0}=\mathbf{u}^{t-1}_{0}+\sum_{i=0}^{j}\Delta\tilde{\mathbf{u}}^{t+i}_{0}$.
3.2 Denoising Diffusion Probalistic Models (DDPMs) for mesh-based simulations
To fully utilize Denoising Diffusion Probalistic Models (DDPMs)’s potential for generating rich, multi-frequency solutions, prediction models must handle multi-scale information. Hierarchical Graph Neural Networks (HGNNs) are particularly well-suited for this, as their architecture inherently learns representations at varying levels of granularity, mirroring the diverse frequency content present in DDPM outputs. Leveraging this idea, we train HGNN to predict the discrete denoising velocity $\mathbf{v}_{i,\theta}(\mathbf{u}^{t}_{i,k},k,\mathbf{u}^{t-1}_{i,0})$ for mesh-based simulations on the mesh graph $\mathcal{G}=(\mathcal{V},\mathcal{E}^{\text{M}})$. It takes the current noisy sample $\mathbf{u}^{t}_{i,k}$ and is conditioned on the previous clean sample $\mathbf{u}^{t-1}_{i,0}$.
Root-node AMG-based Mesh Coarsening. To apply HGNN, we first create a hierarchical mesh graph $\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}})$ consisting of $L+1$ graphs with nodes $\mathcal{V}^{0:L}$ and mesh edges $\mathcal{E}^{0:L,\text{M}}$ using the mesh graph $\mathcal{G}$. We define the fine mesh graph $\mathcal{G}^{0}:=\mathcal{G}$ as our mesh graph at level $l=0$ and coarsen it $L$ times. For coarsening, we use an algebraic root-node-based smooth aggregation solver [53] implemented in [59]. Starting with the adjacency matrix (with self-loops) of the fine mesh $A^{0}$, the solver creates a hierarchy of adjacency matrices $A^{0:L}$.
We use root-node-based smoothed aggregation, which leverages algebraic information from the adjacency matrix, to construct coarse graphs $\mathcal{G}^{1:L}$ and inter-level transfer operators. The selected root nodes $j\in\mathcal{V}^{1:L}$, a subset of the fine mesh nodes, form the coarse graph nodes. We reuse the smoothed aggregation mapping from the AMG algorithm, assigning each fine node $i\in\mathcal{V}^{l}$ to exactly one root node $j\in\mathcal{V}^{l+1}$, to define coarse graph edges. Following [52], we add an edge between $i,j\in\mathcal{V}^{l+1}$ if their aggregation nodes $\mathcal{V}^{l}$ of graph $\mathcal{G}^{l-1}$ are connected. Unlike bi-stride pooling with Delaunay remeshing [22], this AMG-based approach better preserves mesh geometry by leveraging strength connection in the adjacency matrix. Figure 3 visualizes this difference.
We additionally define $L-1$ downsampling edges $\mathcal{E}^{l,\text{D}}$ and upsampling edges $\mathcal{E}^{l,\text{U}}$, which connect the nodes of the graphs $\mathcal{G}^{l}$ and $\mathcal{G}^{l+1}$, and vice versa. We define these as the connections (non-zero values) of the restriction and prolongation matrices of the AMG hierarchy, resulting in an extended hierarchy graph $\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}}\cup\mathcal{E}^{0:L-1,\text{D}}\cup\mathcal{E}^{0:L-1,\text{U}})$. Those matrices are constructed by smoothing the sparse initial aggregation mapping using the adjecency matrix, resulting in a denser matrix with a wider receptive field. As before, we found that the obtained pooling mappings respect the mesh geometry well, as shown in Figure 3 d).
Figure 3: Comparison of AMG-based mesh coarsening to BSDL. a) the original, fine mesh. b) the mesh after two BSDL coarsening steps. c) the mesh after one AMG coarsening steps. This mesh has approximately as many nodes as the mesh in b). d) up- and downsampling edges from level 00 to $1$ (top) and level $1$ to $2$. Bright blue indicates coarse nodes.
For contact experiments, we further extend the graph hierarchy $\mathcal{G}^{0:L}$ by adding contact edges $\mathcal{E}^{0,\text{C}}$ [22] between nodes. In a simulation involving two colliding components, we define a bidirectional edge between the nodes of the first part and the nodes of the second part if their distance is less than the specified contact radius $R$. With these edges, we obtain a hierarchy $\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}}\cup\mathcal{E}^{0:L,\text{C}}\cup\mathcal{E}^{0:L-1,\text{D}}\cup\mathcal{E}^{0:L-1,\text{U}})$.
3.3 Algebraic-hierarchical Message Passing Networks (AMPNs).
Encoder. Let $\mathcal{G}^{0:L}$ be a hierarchical graph as defined above and $\mathbf{u}^{t}_{i,k}$ the noisy sample at denoising step $k$ of simulation step $t$. We define node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{0:L}$, mesh edge embedding $\mathbf{e}_{ij}^{\text{M}}\in\mathcal{E}^{0:L,\text{M}}$, contact edge embeddings $\mathbf{e}_{ij}^{\text{C}}\in\mathcal{E}^{\text{C}}$, downsampling edge embeddings $\mathbf{e}^{\text{D}}_{ij}\in\mathcal{E}^{0:L-1,\text{D}}$ and upsampling edge embeddings $\mathbf{e}^{\text{U}}_{ij}\in\mathcal{E}^{0:L-1,\text{U}}$. We add relative node distances $\mathbf{x}^{t}_{ij}=\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}$ and their norm $|\mathbf{x}^{t}_{ij}|$ to all edge embeddings and the initial distance $\mathbf{x}^{0}_{ij}=\mathbf{x}^{0}_{i}-\mathbf{x}^{0}_{j}$ and their norm $|\mathbf{x}^{0}_{ij}|$ to mesh edges, down- and upsampling embeddings. Node embeddings include a one-hot encoding $n_{i}$ of the node type. Node embeddings at level $l=0$ additionally include $\mathbf{u}^{t}_{i,k}$ and task-specific features. All embeddings are projected to the latent dimension $d$ via linear layers. We add a Fourier encoding [57] for the denoising step $k$ and the normalized level $l^{*}=l/L$ to inform the AMPN of relative graph depth.
Figure 4: Example predictions of ROBIN on the considered datasets. ROBIN predicts the part deformations as well as the von Mises stress (color, yellow is high) on all experiments. a) BendingBeam considers global part deformations of beams induced by local forces. b) In ImpactPlate, the models have to predict locally large deformations, caused by a collision with an accelerated ball. c) The hyperelastic plates in DeformingPlate are deformed by a scripted actuator.
Processor and Decoder. Similar to AMG solvers [54, 53] and UNets [60]Algebraic-hierarchical Message Passing Networks (AMPNs) use a V-cycle to propagate information between levels. They consist of five core message passing modules: Pre-Processing, Downsampling, Solving, Upsampling and Post-Processing, as shown in Figure 1. Pre-Processing, Solving and Post-Processing modules use an  Intra-Level-Message Passing Stack (Intra-MP-Stack) consisting of $N$ message passing steps to update the heterogeneous subgraph $\mathcal{G}^{l}=(\mathcal{V}^{l},\mathcal{E}^{l,\text{C}}\cup\mathcal{E}^{l,\text{M}})$ of level $l$. Given node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{l}$, contact edge embeddings $\mathbf{e}_{ij}^{\text{C}}\in\mathcal{E}^{l,\text{C}}$ and mesh edge embeddings $\mathbf{e}_{ij}^{\text{M}}\in\mathcal{E}^{l,\text{M}}$, the message passing update of the level graph at step $n$ is defined by
$\mathbf{e}_{ij}^{\text{C},n+1}$ $=W^{n}_{\theta,\mathcal{E}^{\text{C}}}~{}\mathbf{e}_{ij}^{\text{C},n}+f^{n}_{\theta,\mathcal{E}^{\text{C}}}(\mathbf{k}^{n}_{i},\mathbf{k}^{n}_{j},\mathbf{e}_{ij}^{\text{C},n})\textrm{ , }$ (2)
$\mathbf{e}_{ij}^{\text{M},n+1}$ $=W^{n}_{\theta,\mathcal{E}^{\text{M}}}~{}\mathbf{e}_{ij}^{\text{C},n}+f^{n}_{\theta,\mathcal{E}^{\text{M}}}(\mathbf{k}^{n}_{i},\mathbf{k}^{n}_{j},\mathbf{e}_{ij}^{\text{M},n})\textrm{ , }$
$\mathbf{k}^{n+1}_{i}$ $=W^{n}_{\theta,\mathcal{V}}~{}\mathbf{k}^{n}_{i}+f^{n}_{\theta,\mathcal{V}}(\mathbf{k}^{n}_{i},\bigoplus_{j}\mathbf{e}_{ij}^{\text{C},n+1},\bigoplus_{j}\mathbf{e}_{ij}^{\text{M},n+1})\quad\text{.}$
The operator $\bigoplus$ denotes a permutation-invariant aggregation, $f^{n}_{\theta,.}$ MLPs and $W^{n}_{\theta,.}$ weight matrices [46, 5, 52]. Downsampling modules update the subgraph $\mathcal{G}^{l,\text{D}}=(\mathcal{V}^{l+1}\cup\mathcal{V}^{l},\mathcal{E}^{l,\text{D}})$ with an  Inter-Level-Message Passing Stack (Inter-MP-Stack) of $N$ message passing steps. The receiver embeddings are $\mathbf{k}^{\text{rec}}_{i}\in\mathcal{V}^{l+1}$, the sender embeddings $\mathbf{k}^{\text{send}}_{j}\in\mathcal{V}^{l}$, and the edge embeddings $\mathbf{e}_{ij}\in\mathcal{E}^{l,\text{D}}$. Similarly, the upsampling layers update the embeddings of the subgraph $\mathcal{G}^{\text{U}}_{l}=(\mathcal{V}^{l+1}\cup\mathcal{V}^{l},\mathcal{E}^{l,\text{U}})$ of level $l$. Here, the receiver embeddings are $\mathbf{k}^{\text{rec}}_{i}\in\mathcal{V}^{l}$, the sender embeddings $\mathbf{k}^{\text{send}}_{j}\in\mathcal{V}^{l+1}$, and the edge embeddings $\mathbf{e}_{ij}\in\mathcal{E}^{l,\text{U}}$. A message passing step of an Intra-MP-Stack is defined as
$\mathbf{e}_{ij}^{n+1}$ $=W^{n}_{\theta,\mathcal{E}}~{}\mathbf{e}_{ij}^{n}+f^{n}_{\theta,\mathcal{E}}(\mathbf{k}^{\text{rec},n}_{i},\mathbf{k}^{\text{send},n}_{j},\mathbf{e}_{ij}^{n})\textrm{ , }$ (3)
$\mathbf{k}^{\text{rec},n+1}_{i}$ $=W^{n}_{\theta,\mathcal{V}}~{}\mathbf{k}^{\text{rec},n}_{i}+f^{n}_{\theta,\mathcal{V}}(\mathbf{k}^{\text{rec},n}_{i},\bigoplus_{j}\mathbf{e}_{ij}^{n+1})\quad\text{.}$
Our V-cycle starts at level $l=0$ with pre-processing and downsampling at each level, repeated until the coarsest level $l=L$ is reached. We then apply multiple message passing steps at level $L$, which has the largest receptive field. Next, we upsample and post-process each level back up to $l=0$. All Pre-Processing, Downsampling, Upsampling, and Post-Processing modules share weights across levels. A final linear layer decodes the fine-level node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{0}$ to produce the velocity prediction $\mathbf{v}_{i,\theta}(\mathbf{u}^{t}_{i,k},k,\mathbf{u}^{t-1}_{i,0})$.
4 Experiments
Figure 5: Left: Rollout error measured by the RMSE of the predicted nodes positions. ROBIN surpasses the accuracy of the baselines on all three datasets BendingBeamImpactPlate and  Deformingplate. Right: Comparison of inference time and error of ROBIN and its variants on BendingBeam. ROBIN achieves the same accuracy as conventional diffusion inference (ROBIN - $m20$), while the inference speed is close the one step variant (ROBIN - OS). Reducing the denoising step $K$, denoted as ROBIN - $K10$ and ROBIN - $K5$, trades accuracy for speed.
Datasets. We evaluate our model on the three different datasets, namely BendingBeam, ImpactPlate [22] and DeformingPlate [5]. We introduce the BendingBeam dataset (Figure 4a)), featuring quasi-static, geometrically non-linear deformations of beams with high aspect ratios. The setup challenges models to capture global deformations via broad receptive fields and resolve high spatial frequencies due to locally thin, low-stiffness regions. In ImpactPlate (Figure 4 b)), the models must learn flexible dynamics with varying material parameters and accurately resolve very localized deformation at the contact point. DeformingPlate (Figure 4 c)) considers quasi-static contact simulations induced by scripted actuators that deform $3$D plates consisting of nonlinear, hyperelastic material.
Experimental setup. For all tasks, we target the displacement residual of the node positions with respect to the next time step $\Delta\mathbf{u}^{t}_{i,0}=\mathbf{x}^{t+1}_{i}-\mathbf{x}^{t}_{i}$ during denoising. We additionally denoise the von Mises stress $\sigma^{t}_{\text{vMises},i,0}$ directly without residuals to gain further insight into the dynamics of the three experiments. ROBIN uses $K=20$ denoising steps and a denoising stride of $m=1$ by default. The task-specific features are specified in Appendix C, listed in Table 2. We measure the prediction error using an RMSE, as specified in Appendix C. Appendix C provides also information about additional settings of ROBIN, including training details and hyperparameters.
Baselines. We compare our model with two prominent baselines for nonlinear deformation simulations, namely MGNs and Hierarchical Contact Mesh Transformes. In Appendix D, we describe the baselines and their setup in more detail.
Variants. We demonstrate that trained ROBINs can easily switch between different rollout modes by varying the denoising stride $m$. We compare ROBIN to conventional autoregressive diffusion inference (ROBIN - $m20$), i.e., $m=20$ denoising steps per time step, and to an intermediate variant (ROBIN - $m5$) with a denoising stride of $m=5$. Additionally, without retraining, we can run ROBIN as one step model, namely ROBIN - 20/OS, which uses the clean sample $\Delta\tilde{\mathbf{u}}^{t_{b}}_{0}\approx\Delta\tilde{\mathbf{u}}^{t_{b}}_{0|K}$ of the first denoising step as prediction. Finally, we train ROBIN with fewer diffusion steps $K=5$ (ROBIN - $K5$) and $K=10$ (ROBIN - $K10$), running both with a denoising stride of $m=1$.
Ablations. We ablate key components of ROBIN to assess their impact. No diffusion trains the AMPN as a one-step autoregressive model with an MSE loss. No hierarchy disables hierarchical message passing, operating only on the fine mesh $G^{0}$. State prediction replaces residual-based prediction with direct denoising of $\mathbf{u}^{t}_{i,k}$ instead of $\Delta\mathbf{u}^{t}_{i,k}$. No shared layer uses $15$ unshared message passing layers. Appendix D provides additional implementation and training details.
5 Results
Baselines. Figure 5 a) comparse the rollout RMSE of ROBIN to HCMT and MGN. ROBIN yields substantial error reductions for the prominent ImpactPlate and DeformingPlate datasets, and achieves an even more remarkable improvement on BendingBeam. Figure 6 demonstrates that ROBIN is able to propagate local boundary conditions accross the part for accurate predicitons of the global part deformation. In addition, ROBIN resolves the non-linear bending curve of the FEM. This requires to combine geometric features at different scales, such as global part dimensions with local thin walls, for an accurate prediction of the global part stiffness.
Figure 6: Comparison of the predicted rollout deformations and von Mises stresses (color, yellow is large) on BendingBeam between a) the FEM, b) ROBIN, c) HCMT, and d) MGN. ROBIN is able to accurately reproduce the FEM results. Both HCMT and MGN are unable to resolve global deformation modes, illustrating the importance of the AMPN for global message propagation.
Inference speed. In Figure 5 b), the rollout RMSE is plotted over the inference time for the BendingBeam dataset. ROBIN is an order of magnitude faster than conventional autoregressive diffusion inference, denoted as ROBIN - $m20$, without compromising accuracy. Most notably, the fastet variant ROBIN - OS, which is a one step inference variant of ROBIN, significantly outperform the accuracy of the baselines (cf. Figure 5 a)). Despite $K=20$ denoising steps, ROBIN requires only $\approx 71\%$ longer than the one step variant, due to the parallel denoising scheme ROBI. By training ROBIN with fewer denoising steps, we obtain models, such as ROBIN - $K10$ and ROBIN - $K5$, whose speed and accuracy lies between the one-step variant and ROBIN. We obtain similar results for ImpactPlate and DeformingPlate, which is visualized in Figure 9 in Appendix E.
Ablations. As visualized in Figure 7, the ablation results demonstrate the effectiveness of the individual components of ROBIN. Non-shared layers decrease the performance significantly on BendingBeam, due the the broader receptive field. The large error increase of the State prediction ablations indicates that residual prediction are crucial for Lagrangian simulations. ROBIN outperforms the non-hierarchical architecture variant as well as the non-diffusion variant on all datasets, demonstrating the synergy between DDPMs and AMPNs.
Figure 7: Rollout RMSE of the displacement predictions. Shared layers are crucial for BendingBeam to increase the receptive field. Replacing residual predictions with state predictions, hierarchal architectures with non-hierarchical architectures, and diffusion with non-diffusion architectures significantly decrease the accuracy across all datasets.
6 Conclusion
We introduced Rolling Diffusion-Batched Inference Network (ROBIN), a diffusion-based HGNN that utilizes AMPNs to refine mesh-based predictions across scales. Leveraging the expressiveness of multiscale message passing and the accuracy of diffusion, ROBIN outperforms state-of-the-art simulators on varied nonlinear solid mechanics tasks in terms of predictive accuracy. These tasks include a novel BendingBeam dataset that reveals limitations of current learned simulators. ROBI, ROBIN’s inference scheme, parallelizes diffusion across time steps, reducing inference runtime by up to an order of magnitude without sacrificing accuracy. We validated ROBIN on three challenging datasets, including the new BendingBeam benchmark, and demonstrated significant gains in accuracy and efficiency. We discuss the broader impact of this work and our method in Appendix A.
Limitations and Future Work. ROBIN currently does not possess SO(3) equivariance, limiting its applicability to predictions involving tensor fields or anisotropic materials. We focus on DDPM, while other diffusion formulations or denoising schedules could provide valuable insights and further enhance performance. Similarly, our experiments cover nonlinear solid mechanics, and extensions to other domains, such as fluid dynamics, are a promising direction. Lastly, ROBIN’s combination of fast inference and high accuracy opens opportunities for accelerating multi-stage design and optimization workflows.
Acknowledgements
This work is part of the DFG AI Resarch Unit 5339 regarding the combination of physics-based simulation with AI-based methodologies for the fast maturation of manufacturing processes. The financial support by German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) is gratefully acknowledged. The authors acknowledge support by the state of Baden-Württemberg through bwHPC, as well as the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the German Federal Ministry of Education and Research. This work is supported by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT partition.
Appendix A Broader Impact
The ML-based simulator,  Rolling Diffusion-Batched Inference Network (ROBIN), offers significant advantages for computational modeling and simulation. This is achieved by significantly reducing computational costs while producing accurate simulations. This enables engineers to iterate through significantly more design variations or to quickly evaluate numerous scenarios using the fast model. However, like all powerful computational tools, there is a risk of misuse, for instance, in weapons development or unsustainable resource exploitation.
Appendix B Datasets
Table 1 provides on overview over the considered datasets in this work.
Table 1: Comparison of the three datasets considered in this work. The column Nonlinearity distinguishes three different types: geometry (Geo), material (Mat), and boundary conditions (BC).
Datasets Dynamic Nonlinearity Simulator Mesh Type Steps T Dim
BendingBeam Quasi-Static Geo Scikit-fem Triangles 400 2D
ImpactPlate Dynamic Geo, BC ANSYS Triangles 52 2D
DeformingPlate Quasi-Static Geo, BC, Mat COMSOL Tetrahedrons 400 3D
BendingBeam. The dataset BendingBeam considers the bending of beam parts due to external forces. Bending is one of the most basic deformation modes of parts in structural mechanics. The dataset is designed as a diagnostic benchmark for neural PDE solvers, addressing various potential bottlenecks. The force and handle boundary conditions are very local, only beeing defined on a small subset of mesh nodes (cf. Figure 8). However, the resulting deformations affect all nodes.
Figure 8: Visualization of the different node types on the BendingBeam experiment. Blue nodes handle boundaries with fixed node positions. The red color indicates nodes where the force boundary condition is applied. Those locally applied forces in combination with the global, geometry-dependent part stiffness mainly determine the global deformation.
Hence, the neural PDE must be very good at propagating local information to all nodes of the mesh. Next, the dataset considered beams with large aspect ratios. This results in large graph diameters, which represent the shortest path between the most distant nodes. The mesh resolution is increased at thin walls, which additionally increases the graph diameter. The local boundary conditions has to be transmitted across a large number of nodes, which challenges the ability to propagate messages globally. The geometry and especially the thin locations of the geometry strongly influence the global bending stiffness and deformation of the part. Overall, the model has to output accurate solutions across various spatial frequencies.
The solutions are created with scikit-fem [61], iteratively solved using Netwon-Raphson until the residual dropped below a tolerance of $10^{-8}$. Each simulation is solved for a total number of $400$ time steps. We create a total number of $1000$ simulations for training, $100$ for validation and $100$ for testing.
Appendix C Setup
Hardware and Compute. We train all models on a single NVIDIA A100 GPU with a maximum training time of $48$ hours, while most models required approximately $40$ hours. In total, we trained on $3$ datasets $9$ models each on $5$ seeds: ROBINMGNHCMT, as well as $4$ ablations and $2$ variants of ROBIN. That amounts to $40~{}\text{hours}\times 5\times 3\times 8=5400~{}\text{hours}$ training time. We required a comparable amount of time for development and hyperparameter tuning. In addition, we run inference experiments to measure the inference speed of $6$ model variants each trained on $5$ different seeds. We run the inference experiments on $3$ datasets with approximately a mean of $1$ hour per experiment. In total we obtain a runtime of $1~{}\text{hour}\times 5\times 3\times 6=90~{}\text{hours}$ for the inference experiments.
Training. We implement ROBIN in Pytorch [62] and train it with ADAM [63]. We use an exponential learning rate decay, which decrease the learning rate from $41e-$ 4 to $61e-$ 6 over the training time, including $1000$ linear increasing warm up steps. We clip gradients such that their $L_{2}$-norm doesn’t exceed $1$. We train ROBIN in BendingBeam with $9$M samples and in ImpactPlate with $6$M samples both with a batch size of $16$, resulting in $562\text{,}500$ and $375\text{,}000$ training iterations. In DeformingPlate we reduce the batch size to $12$ and train for $300\text{,}000$ iterations with $3.6$M samples.
Features. Table 2 provides an overview over the used input and ouput features for ROBIN. In addition to the default features, we extend the node embeddings of BendingBeam with the force residual $\Delta f_{\text{BC}}$, which is defined by the boundary condition. In ImpactPlate, we add the density $\rho_{i}$ and the Young’s modulus $Y_{i}$ as node features. In DeformingPlate, we add the scripted displacement residual $\Delta\mathbf{u}_{\text{BC}}$ of the actuator. We normalize all input features based on the training dataset, setting them to have a zero mean and unit variance. We add a small amount of training noise [4, 5] of $10^{-5}~{}\mathbf{\sigma}_{\mathbf{x}}$ to the node positions $\mathbf{x}_{i}^{t}$, where we scale the noise level with the standard deviation of the features $\mathbf{\sigma}_{\mathbf{x}}$. For ImpactPlate we noise the input history $\Delta\mathbf{u}^{t-1}_{i,0}=\mathbf{x}^{t}_{i}-\mathbf{x}^{t-1}_{i}$ with $10^{-3}~{}\sigma_{\mathbf{x}}$ to prevent overfitting on the history.
Table 2: Node $\mathbf{k}_{i}$ and edge embeddings $\mathbf{e}_{ij}$ for the different datasets, depending on the node $\mathcal{V}$ and edge sets $\mathcal{E}$.
Datasets Inputs $\mathcal{V}^{0}$ Inputs $\mathcal{V}^{1:L}$ Inputs $\mathcal{E}^{0:L,\text{M}}$ Inputs $\mathcal{E}^{0:L,\text{C}}$ Inputs $\mathcal{E}^{0:L,\text{U/D}}$ Outputs $\mathcal{V}^{0,\text{M}}$
BendingBeam $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta f_{\text{BC}}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
ImpactPlate $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta\mathbf{u}^{t-1}_{i,0}$, $\rho_{i}$, $Y_{i}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
DeformingPlate $\mathbf{n}_{i}$, $\Delta\mathbf{u}^{t}_{i,k}$, $\Delta\mathbf{u}_{\text{BC}}$ $\mathbf{n}_{i}$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$ ${\mathbf{x}}^{t}_{ij}$,$|{\mathbf{x}}^{t}_{ij}|$, ${\mathbf{x}}^{0}_{ij}$,$|{\mathbf{x}}^{0}_{ij}|$ ${\mathbf{v}}_{i,\theta}(\Delta\mathbf{u}^{t}_{i,k})$
Hierarchical Graph. Since the relative motion of the components in the considered experiments is not too large, we define the contact edges based on the initial mesh configuration and keep them constant to maintain a constant graph. In DeforminPlate we set the contact radius to $R=0.1$, connecting actuator nodes with plate nodes. In ImpactPlate we connect ball nodes and plate nodes with a radius of $R=1.2$. In all three experiments, we create $L=2$ coarse layers to obtain $3$ mesh levels.
Algebraic-hierarchical Message Passing Networks. We use $3$ Pre- and $3$ Post-processing layers, $2$ Up- and $2$ Downsampling layers and $5$ Solving layers, which yields a total number of $15$ learnable layers. We add a layer norm before each MLP and use two linear layers, a hidden size of $128$ and a  Sigmoid Linear Unit (SiLU[64] activation function. A max aggregation is used in all message passing layer.
Denoising Diffusion Probalistic Models. We use $K=20$ denoising steps and a denoising stride of $m=5$ for ROBIN by default. The $\beta$ variances of the DDPM scheduler are geometrically spaced for training and inference, starting from a minimum noise variance of $41e-$ 4 (for $\beta_{1}$) and going up to $1.0$ (for $\beta_{K}$).
Metrics. To compare the rollout accuracy, we follow [22] and define the Root Mean Squared Euclidean distance error $\text{RMSE}=\sqrt{1/N_{i}\sum_{i=1}^{N_{i}}\sum_{j=1}^{N_{j}}(\tilde{u}_{ij}-u_{ij})^{2}}$ between the prediction $\tilde{u}_{ij}$ and the ground truth ${u}_{ij}$ both with $N_{i}$ nodes and $N_{j}$ features. We then calculate the mean over all time steps, the mean over the dataset and finally the mean $\mu$ and standard deviation $\sigma$ over the $5$ seeds.
Appendix D Baselines, Ablations and Variants
Baselines. We use the official TensorFlow [65] implementation of the authors for the baselines HCMT [^2] https://github.com/yuyudeep/hcmt/tree/main [22] and MGN [^3] https://github.com/google-deepmind/deepmind-research/tree/master/meshgraphnets [5]. We use ADAM [63] for training HCMT and MGN with an exponential leanring rate deacy form $41e-$ 4 to $61e-$ 6, a batch size of $1$ and a hidden size of $128$. We use on all datasets a total number of $15$ message passing steps for MGN, as well as a total of $15$  Hierarchical Mesh Transformer (HMT) and  Contact Mesh Transformer (CMT) layers for HCMT.
On BendingBeam, we train HCMT for $4$M training iterations, use a contact radius of $R=0.01$ and a training noise [4, 5] of $0.001$. We maximize the receptive field and set the number of mesh levels to $5$, the maximum at which at least five nodes remain available across all meshes in the dataset, required for Delaunay remeshing [66]. This results in $9$ HMTs. Since BendingBeam is a contact-free task, we replaced the dual-branch CMT by $6$ single-branch Mesh Transformer layers, that only attends to mesh edges instead of mesh and contact edges. We use the same architecture and hyperparameter for HCMT on ImpactPlate and DeformingPlate as proposed by the authors [22], and train it for $2$M steps and $3$M steps, respectively.
We follow the authors implementation and add world edges to the mesh graph of MGN instead of contact edges to increase the receptive field of the non-hierarchical architecture. MGN is trained for $3$M iterations on BendingBeam and uses a training noise of $0.001$ with a world edge radius of $R=0.13$. On ImpactPlate we train MGN for $3$M steps and use a world edge radius of $R=0.03$ and a training noise of $0.003$. We train MGN on DeformingPlate for $1.5$M steps and use the authors’ proposed settings [5]. To prevent out-of-memory errors in edge cases on DeformingPlate, we restrict the number of world edges to $200\text{,}000$ by selecting those with the smallest node distances.
Ablations. For the No hierarchy ablation we use the fine mesh graph $\mathcal{G}^{0}$ instead of the hierarchical graph $\mathcal{G}^{0:L}$, replace our AMPN by a single Intra-MP-Stack with $15$ learnable message passing steps and remove the positional level encoding. In addition, we follow MGN and replace contact edges by world edges to increase the receptive field. To stay within the training budget, we reduce the number of trainings samples to 1.21.2{1.2}1.2M for BendingBeam and to 88{8}8M for ImpactPlate. For DeformingPlate we reduce the batch size to $1$ and the training samples to 0.60.6{0.6}0.6M and also restrict the number of world edges to $200\text{,}000$ as for MGN. The No diffusion ablation trains the AMPN with an MSE loss to predict directly the displacement residual $\Delta u^{t}$ and use an one step autoregressive rollout, such as HCMT and MGN. We use the same training noise settings as the baselines to stabilize the rollouts. The No shared layer ablation uses a total number of $15$ non-shared learnable message passing layers, that are distributed as follows: $1$ Pre-Processing and $1$ Post-Processing layer per level, $1$ Up- and $1$ Downsampling layer between each level, and $5$ Solving layers. The faster predictions allows us to increase the number of training samples to $1${11}11M for BendingBeam, to 88{8}8M for ImpactPlate, and to 4.64.6{4.6}4.6M for DeformingPlate.
Appendix E Results
Figure 9 visualize the displacement rollout RMSE over the inference time of on ImpactPlate and DeformingPlateROBIN is compared to different inference modes and trained variants with fewer diffusion steps.
(a) ImpactPlate.
(b) DeformingPlate.
Figure 9: RMSE rollout error and inference time of ROBIN, different inference modes of ROBIN, and variants with fewer denoising steps $K$ on a) ImpactPlate and b) DeformingPlateROBIN is most accurate on ImpactPlate and on par with the slower variants ROBIN - $m5$ and ROBIN - $m20$. The one step variant ROBIN - OS achieves the largest speed-up on DeforminPlate, but also loses the most accuracy there. Reducing the diffusion steps to $10$ (ROBIN - $K10$) and $5$ (ROBIN - $K5$) denoising steps, respectively, increases speed while decreases the accuracy.
Figure 10 and Figure 11 visualize the rollout displacement and von Mises stress prediction of ROBINHCMT and MGN on ImpactPlate and DeformingPlate, respectively.
Figure 10: Comparison of the rollout deformation prediction and von Mises stress prediction (color) on ImpactPlate to the ground truth of the FEMROBIN most accurately resolves the deformation at the contact surface and the resulting stress. The deformation prediction of HCMT comes close to the FEM prediction while the stress is underestimated. MGNs predicts accurately the global modes but exhibits local disturbances.
Figure 11: Rollout deformation and von Mises stress prediction (color) on DeformingPlate of ROBIN, the baselines and the FEM. All models accurately reproduce the part deformation. HCMT overestimates slightly the stress at the thin wall between the hole and the boundary.